OpenStudy (anonymous):

Identify the 8th term of the geometric sequence -2, -12, -72 ... a.-3,359,232 b. -559,872 c. 559,872 d. 3,359,232

Parth (parthkohli):

\[{\rm A}_n = {\rm A} \times r^{n - 1}\]

OpenStudy (anonymous):

A= first term =-2 r=Geometric ratio = -12/(-2) =6

OpenStudy (anonymous):

Now, find A_8

Parth (parthkohli):

One may notice that \({\rm A} = -2\) and \(r = \dfrac{\rm A_2}{\rm A_1} = 6\)

OpenStudy (anonymous):

ive already tried solving it myself it just wasnt working out

Parth (parthkohli):

\[A _8 = -2 \times 6^{8 - 1}\]

OpenStudy (anonymous):

is the answer b?